Jason has a problem. As a graduation gift, his father offered him either an ₱80,000 cash today or an investment that will give him ₱100,000 two years in the future. Assuming a prevailing interest rate of 12%, which option should Jason choose to maximize the money he can receive?

Answer by Theo(13342)

he can get 80,000 today or 100,000 two years from now.
the prevailing interest rate he can earn on the money he receives today is 12% per year.
the formula to use is f = p * (1 + r) ^ n
f is the future value
p is the present value
r is the interest rate per time period
n is the number of time periods.
use this formula to see how much he would have in two years if he took the 80,000 today.
the formula becomes f = 80,000 * (1 + .12) ^ 2 = 100,352.
he should take the 80,000 today and invest it at 12%.
he'll have more than 100,000 in two years, but not by much.
still, it's more, so that's the way to go.
if the money he invests today is invested at 12% compounded monthly, then he'll have even more.
with monthly compounding, the formula becomes f = 80,000 * (1 + .12/12) ^ (2 * 12) = 101,578.7719.
that makes taking the 80,000 today and investing it at 12% more compelling.